# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 5. Lifting to $p$-adic groups
In this section, we give an application of Theorem 4.4 to the depth zero case of irreducible smooth representations of ${\rm GL}_{n}(F)$, where $F$ is a non-Archimedean local field.
@@ 5.1. Local base change lifting
For a non-Archimedean local field $K$, let $\mathcal{A}(n,K)$ be the set of isomorphism classes of irreducible $\ell$-adic smooth representations of ${\rm GL}_{n}(K)$. Let $\mathcal{G}(n,K)$ be the set of isomorphism classes of semi-simple $\ell$-adic Weil–Deligne representations of the Weil group $\mathcal{W}_{K}$. The $\ell$-adic local Langlands correspondence is a bijection
@@
which is uniquely characterized by certain identities of local constants that one attaches to the elements of $\mathcal{A}(n,K)$ and $\mathcal{G}(n,K)$ (see [HT01], [Hen00], [Sch13]). Moreover, the set of irreducible $\ell$-adic cuspidal representations of ${\rm GL}_{n}(K)$ is mapped onto the set of $n$-dimensional irreducible $\ell$-adic representations of $\mathcal{W}_{K}$ via ${\rm LLC}_{K}$. Note that the classical local Langlands correspondence is a bijection between the isomorphism classes of irreducible smooth complex representations of ${\rm GL}_{n}(K)$ and the isomorphism classes of $n$-dimensional complex semisimple Weil–Deligne representations. To get a correspondence over $\overline{\mathbb{Q}}_{\ell}$, one twists the original correspondence by the character $\nu^{(1-n)/2}$, where $\nu$ is the composition of the determinant character of ${\rm GL}_{n}(K)$ and the absolute value of $K$. For details, see [Clo90, Conjecture 4.4, Section 4.2] and [Hen01, Section 7].
Say $F$ be a non-Archimedean local field and let $E$ be a Galois extension of $F$ such that $[E:F]=\ell$. Let $\pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(F)$. Let $\Pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(E)$ such that
@@
5.1.1.
Let $K_{0}(F)$ be the compact open subgroup ${\rm GL}_{n}(\mathfrak{o}_{F})$, and let $K_{1}(F)$ be the pro-$p$ subgroup
@@
where $M_{n}(\mathfrak{o}_{F})$ is the ring of $n\times n$ matrices with entries in the ring of integers $\mathfrak{o}_{F}$, and $\varpi_{F}$ is a uniformiser of $F$. The quotient group $K_{0}(F)/K_{1}(F)$ is isomorphic to ${\rm GL}_{n}(\mathbb{F}_{q})$, where $\mathbb{F}_{q}$ is the residue field of $F$. An irreducible smooth representation $(\pi,V)$ of ${\rm GL}_{n}(F)$, where $V$ is a vector space over $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$, is said be of depth-zero if there exists a non-zero vector in $V$ that is fixed under the action of $K_{1}(F)$.
5.1.2.
In this part, we discuss the compatibility of the local base change and Shintani lift. Let $\mathcal{A}(n,F)_{0}$ be the set of isomorphism classes of irreducible smooth depth-zero $\ell$-adic representations of ${\rm GL}_{n}(F)$, and let $\mathcal{A}(n,\mathbb{F}_{q})$ be the set of isomorphism classes of irreducible $\ell$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. In [SZ99, Section A.1.3], Schneider–Zink defined a map
@@
Let $E$ be a finite unramified Galois extension of $F$ with $[E:F]=\ell$. In [SZ08, Appendix A, equation (A.7)], Zink–Silberger proved the following identity
@@
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$ and $\sigma$ be a generator of ${\rm Gal}(E/F)$. Let $(\Pi,V)$ be an absolutely irreducible depth-zero integral generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\simeq\Pi^{\sigma}$. Let $(\pi,W)$ be the irreducible depth-zero $\ell$-adic integral cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. From the definition, we have $\varphi_{0}^{E}(\Pi)=\Pi^{K_{1}(E)}$ and $\varphi_{0}^{F}(\pi)=\pi^{K_{1}(F)}$. Using the identity (5.1), we get that the irreducible ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$. Using these observations, we now prove the following theorem.
Theorem 5.1.
@@
    For any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ invariant $\Lambda$-lattice $\mathcal{L}$ of $\Pi$, the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
Let $\mathcal{V}$ denote either the $\Lambda[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $\mathcal{L}$ or the $\overline{\mathbb{F}}_{\ell}[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $r_{\ell}(\Pi)$. Consider the compact open subgroup $K_{1}(E)$ of ${\rm GL}_{n}(E)$. Then we have
$\mathcal{V}=\mathcal{V}^{K_{1}(E)}\bigoplus\mathcal{V}(K_{1}(E)).$
Using the compatibility of Tate cohomology with finite direct sums, we get
@@
Since the ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$, it follows from Theorem 4.4 that $\widehat{H}^{0}(\mathcal{V}^{K_{1}(E)})$ is ${\rm GL}_{n}(\mathbb{F}_{q})$-isomorphic to $r_{\ell}(\pi^{K_{1}(F)})^{(\ell)}$. Since $\widehat{H}^{0}(\mathcal{V})$ is irreducible and cuspidal (Theorem 3.6 and Remark 3.7), and the space $\widehat{H}^{0}(\mathcal{V})^{K_{1}(F)}$ contains $(r_{\ell}(\pi)^{(\ell)})^{K_{1}(F)}$, we get that $\widehat{H}^{0}(\mathcal{V})$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$. This completes the proof. ∎
@@ 5.2. Tate cohomology of depth-zero generic representations
This part aims to prove Treumann–Venkatesh’s conjecture for depth-zero $\ell$-adic integral generic representations of ${\rm GL}_{n}(F)$ without the assumption on the prime $\ell$ made in [ND24], i.e., $\ell$ does not divide $\lvert{\rm GL}_{n-1}(\mathbb{F}_{q})\rvert$ for $n\geq 3$. Let us first introduce some notations.
@@ 5.2.1. Supercuspidal support
Let $\lambda=(n_{1},\dots,n_{r})$ be a partition of $n$. Let $P_{\lambda}$ be the parabolic subgroup of ${\rm GL}_{n}$ consisting of block upper triangular matrices of type $\lambda$. Let $M_{\lambda}$ be the Levi subgroup of $P_{\lambda}$ and is isomorphic to ${\rm GL}_{n_{1}}\times\cdots\times{\rm GL}_{n_{r}}$. Let $R$ be either $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$. For each $1\leq i\leq r$, let $\xi_{i}$ be a smooth irreducible supercuspidal $R$-representation of ${\rm GL}_{n_{i}}(F)$. Let $\xi$ be the $M_{\lambda}(F)$ representation $\xi_{1}\otimes\cdots\otimes\xi_{r}$. Let $\pi$ be a smooth irreducible $R$-representation of ${\rm GL}_{n}(F)$. There exists a pair $(M_{\lambda}(F),\xi)$ as defined above such that $\pi$ is a sub-quotient of the normalised parabolic induction $i_{P_{\lambda}(F)}^{{\rm GL}_{n}(F)}\xi$. Moreover, the pair $(M_{\lambda}(F),\xi)$ is unique upto ${\rm GL}_{n}(F)$-conjugacy and is called the supercuspidal support of $\pi$. When $R=\overline{\mathbb{Q}}_{\ell}$, this follows from the Frobenius reciprocity (see [BZ77] for instance). When $R=\overline{\mathbb{F}}_{\ell}$, this is a deep result, proved by Vignéras ([Vig98]).
5.2.2.
Let $\pi$ be an integral generic $\ell$-adic representation of ${\rm GL}_{n}(F)$, and let $(M(F),\xi)$ be the supercuspidal support of $\pi$. We assume that $\pi$ is a sub-representation of $i_{P(F)}^{{\rm GL}_{n}(F)}\xi$, where $P$ is a parabolic subgroup of ${\rm GL}_{n}$ whose Levi component is $M$. Let $\Pi$ be an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\pi$. Then $\Pi$ is a sub-representation of $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where $\eta_{i}$ is an absolutely irreducible integral generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ for each $1\leq i\leq r$, and $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$.
5.2.3.
For each $1\leq i\leq r$, let $\mathcal{L}_{i}$ be a ${\rm GL}_{n_{i}}(E)\rtimes\Gamma$ stable $\Lambda$-lattice in $\eta_{i}$. Let $\mathcal{L}$ be the module $\mathcal{L}_{1}\otimes\cdots\otimes\mathcal{L}_{r}$, which is a $\Lambda[M(E)\rtimes\Gamma]$ stable lattice in $\eta$. By [Vig96, Chapter I, Section 9.3], $i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$ is a ${\rm GL}_{n}(E)$ stable $\Lambda$-lattice in $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$–which is also stable under the action of $\Gamma$ on $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$, defined by
@@
Theorem 5.2.
Let $\pi$ be a smooth, depth-zero, integral, generic $\ell$-adic representation of ${\rm GL}_{n}(F)$, with $J_{\ell}(\pi)$ the unique generic sub-quotient of the $-\ell$ reduction of $\pi$. Let $\Pi$ be an absolutely irreducible, depth-zero, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. Then the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is the unique generic sub-quotient of the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L}^{\prime})$.
@@
and each $\xi_{i}$ is an irreducible, integral, cuspidal $\ell$-adic representation of ${\rm GL}_{n_{i}}(F)$. Now, $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$, being the base change lift of $\pi$, is a subrepresentation of $i_{P(E)}^{G_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where each $\eta_{i}$ is an irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ such that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$. Then we have the following commutative diagram
@@
where $\varphi$ is a ${\rm GL}_{n}(F)$ equivariant map induced from the injection $\mathcal{L^{\prime}}\hookrightarrow i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$. Note that the twisted Jacquet modules $\mathcal{L}^{\prime}_{N_{n}(E),\Theta_{E}}$ and $(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})_{N_{n}(E),\Theta_{E}}$ are free $\Lambda$-modules of rank $1$. Also note that the vertical maps $f_{1}$ and $f_{2}$ are non-zero, and in fact are isomorphisms. Hence, $\varphi$ takes the generic sub-quotients of $\widehat{H}^{0}(\mathcal{L}^{\prime})$ to the generic sub-quotients of $\widehat{H}^{0}(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})$. Recall that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$, and $\mathcal{L}_{i}$ is a ${\rm GL}_{n_{i}}(E)\rtimes\langle\sigma\rangle$ stable $\Lambda$-lattice in $\eta_{i}$, for each $1\leq i\leq r$. Using Theorem 3.6, we get that the Tate cohomology group $\widehat{H}^{0}(\mathcal{L}_{i})$ is isomorphic to $r_{\ell}(\xi_{i})^{(\ell)}$. Now, using the isomorphism (2.1), we get
@@
Thus, the supercuspidal support of the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is same as the supercuspidal support of the generic sub-quotients of the Tate cohomology $\widehat{H}^{0}(\mathcal{L}^{\prime})$. Hence, $J_{\ell}(\pi)^{(\ell)}$ is a sub-quotient of $\widehat{H}^{0}(\mathcal{L}^{\prime})$. This completes the proof. ∎

### v2

@@ 6. Lifting to $p$-adic groups
In this section, we give an application of Theorem 5.4.3 to the depth zero case of irreducible smooth representations of ${\rm GL}_{n}(F)$, where $F$ is a non-Archimedean local field.
@@ 6.1. Local base change lifting
For a non-Archimedean local field $K$, let $\mathcal{A}(n,K)$ be the set of isomorphism classes of irreducible $\ell$-adic smooth representations of ${\rm GL}_{n}(K)$. Let $\mathcal{G}(n,K)$ be the set of isomorphism classes of semisimple $n$-dimensional $\ell$-adic Weil–Deligne representations of the Weil group $\mathcal{W}_{K}$. The $\ell$-adic local Langlands correspondence is a bijection
@@
which is uniquely characterized by certain identities of local constants that one attaches to the elements of $\mathcal{A}(n,K)$ and $\mathcal{G}(n,K)$ (see [HT01], [Hen00], [Sch13]). Moreover, the set of irreducible $\ell$-adic cuspidal representations of ${\rm GL}_{n}(K)$ are mapped onto the set of $n$-dimensional irreducible $\ell$-adic representations of $\mathcal{W}_{K}$ via ${\rm LLC}_{K}$. Note that the classical local Langlands correspondence is a bijection between the set of isomorphism classes of irreducible smooth complex representations of ${\rm GL}_{n}(K)$ and the set of isomorphism classes of $n$-dimensional complex semisimple Weil–Deligne representations. To get a correspondence over $\overline{\mathbb{Q}}_{\ell}$, one twists the original correspondence by the character $\nu^{(1-n)/2}$, where $\nu$ is the composition of the determinant character of ${\rm GL}_{n}(K)$ and the absolute value of $K$. For details, see [Clo90, Conjecture 4.4, Section 4.2] and [Hen01, Section 7].
Say $F$ be a non-Archimedean local field, and let $E/F$ be a finite Galois extension such that $[E:F]=\ell$. Let $\pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(F)$. Let $\Pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(E)$ such that
@@
6.2.
In this part, we discuss the compatibility of the local base change and Shintani lift. Let $K_{0}(F)$ be the compact open subgroup ${\rm GL}_{n}(\mathfrak{o}_{F})$, and let $K_{1}(F)$ be the $p$ subgroup
@@
where $M_{n}(\mathfrak{o}_{F})$ is the ring of $n\times n$ matrices with entries in the ring of integers $\mathfrak{o}_{F}$, and $\varpi_{F}$ is a uniformiser of $F$. The quotient group $K_{0}(F)/K_{1}(F)$ is isomorphic to ${\rm GL}_{n}(\mathbb{F}_{q})$, where $\mathbb{F}_{q}$ is the residue field of $F$. An irreducible smooth representation $(\pi,V)$ of ${\rm GL}_{n}(F)$, where $V$ is a vector space over $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$, is said be of depth-zero if there exists a non-zero vector in $V$ that is fixed under the action of $K_{1}(F)$. Let $\mathcal{A}(n,F)_{0}$ be the set of isomorphism classes of irreducible smooth depth-zero $\ell$-adic representations of ${\rm GL}_{n}(F)$, and let $\mathcal{A}(n,\mathbb{F}_{q})$ be the set of isomorphism classes of irreducible $\ell$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. In the article [SZ99, Section A.1.3], Schneider–Zink defined a map
@@
Let $E$ be a finite unramified extension of $F$ with $[E:F]=\ell$. In [SZ08, Appendix A, equation (A.7)], Zink–Silberger proved the following identity
@@
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$ and $\sigma$ be a generator of ${\rm Gal}(E/F)$. Let $(\Pi,V)$ be an absolutely irreducible depth-zero integral generic $\mathcal{K}$ representation of ${\rm GL}_{n}(E)$ such that $\Pi\simeq\Pi^{\sigma}$. Let $(\pi,W)$ be the irreducible depth-zero integral $\ell$-adic cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. From the definition, we have $\varphi_{0}^{E}(\Pi)=\Pi^{K_{1}(E)}$ and $\varphi_{0}^{F}(\pi)=\pi^{K_{1}(F)}$. Using the identity (6.1), we get that the irreducible ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible $\ell$-adic cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$. Using these observations, we now prove the following theorem.
Theorem 6.2.1.
@@
    For any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice $\mathcal{L}$ of $\Pi$, the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
Let $\mathcal{V}$ be either the $\Lambda[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $\mathcal{L}$ or the $\overline{\mathbb{F}}_{\ell}[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $r_{\ell}(\Pi)$. Consider the compact open subgroup $K_{1}(E)$ of ${\rm GL}_{n}(E)$. We have the following decomposition
$\mathcal{V}=\mathcal{V}^{K_{1}(E)}\bigoplus\mathcal{V}(K_{1}(E)),$
where $\mathcal{V}(K_{1}(E))=\oplus\mathcal{V}_{\rho}$, where $\rho$ is a non-trivial representation of $K_{1}(E)$ and $\mathcal{V}_{\rho}$ is the $\rho$ isotypic component in $\mathcal{V}$. The space $\mathcal{V}(K_{1}(E))$ is stable under the action of ${\rm Gal}(E/F)$. Using the compatibility of Tate cohomology with finite direct sums, we get
@@
Note that the ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$. Then it follows from Theorem 5.4.3 that $\widehat{H}^{0}(\mathcal{V}^{K_{1}(E)})$ is ${\rm GL}_{n}(\mathbb{F}_{q})$-isomorphic to $r_{\ell}(\pi^{K_{1}(F)})^{(\ell)}$. Since $\widehat{H}^{0}(\mathcal{V})$ is irreducible and cuspidal (Theorem 4.1.1 and Remark 4.1.2), and the space $\widehat{H}^{0}(\mathcal{V})^{K_{1}(F)}$ contains $(r_{\ell}(\pi)^{(\ell)})^{K_{1}(F)}$, we get that $\widehat{H}^{0}(\mathcal{V})$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$. This completes the proof. ∎
@@ 6.3. Tate cohomology of depth-zero generic representations
This part aims to prove Treumann–Venkatesh’s conjecture for $\ell$-adic integral, depth-zero, generic representations of ${\rm GL}_{n}(F)$ without the assumption on the prime $\ell$ made in [ND24], i.e., $\ell$ does not divide $\lvert{\rm GL}_{n-1}(\mathbb{F}_{q})\rvert$ for $n\geq 3$. We first introduce some notations.
@@ 6.3.1. Supercuspidal support
Let $\lambda=(n_{1},\dots,n_{r})$ be a partition of $n$. Let $P_{\lambda}$ be the parabolic subgroup of ${\rm GL}_{n}$ consisting of block upper triangular matrices of type $\lambda$. Let $M_{\lambda}$ be the Levi subgroup of $P_{\lambda}$ which is isomorphic to ${\rm GL}_{n_{1}}\times\cdots\times{\rm GL}_{n_{r}}$. For each $1\leq i\leq r$, let $\xi_{i}$ be an irreducible smooth supercuspidal $R$-representation of ${\rm GL}_{n_{i}}(F)$, where $R$ denotes either $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$. Let $\xi$ be the $M_{\lambda}(F)$ representation $\xi_{1}\otimes\cdots\otimes\xi_{r}$. Let $\pi$ be a smooth irreducible $R$-representation of ${\rm GL}_{n}(F)$. There exists a pair $(M_{\lambda}(F),\xi)$ as defined above such that $\pi$ is a sub-quotient of the normalised parabolic induction $i_{P_{\lambda}(F)}^{{\rm GL}_{n}(F)}\xi$. Moreover, the pair $(M_{\lambda}(F),\xi)$ is unique up to ${\rm GL}_{n}(F)$-conjugacy and is called the supercuspidal support of $\pi$. When $R=\overline{\mathbb{Q}}_{\ell}$, this follows from the Frobenius reciprocity (see [BZ77] for instance). When $R=\overline{\mathbb{F}}_{\ell}$, this is a deep result, proved by Vignéras ([Vig98]).
6.3.2.
Let $\pi$ be an $\ell$-adic integral generic representation of ${\rm GL}_{n}(F)$, and let $(M(F),\xi)$ be the supercuspidal support of $\pi$. We assume that $\pi$ is a subrepresentation of $i_{P(F)}^{{\rm GL}_{n}(F)}\xi$, where $P$ is a parabolic subgroup of ${\rm GL}_{n}$ whose Levi component is $M$. Let $\Pi$ be an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. Then $\Pi$ is a subrepresentation of $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where $\eta_{i}$ is an absolutely irreducible integral generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ and $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$ for each $1\leq i\leq r$.
6.3.3.
For each $1\leq i\leq r$, let $\mathcal{L}_{i}$ be a $\Lambda[{\rm GL}_{n_{i}}(E)\rtimes\Gamma]$ stable lattice in $\eta_{i}$. Let $\mathcal{L}$ be the module $\mathcal{L}_{1}\otimes\cdots\otimes\mathcal{L}_{r}$, which is a $M(E)\rtimes\Gamma$ stable lattice in $\eta$. By [Vig96, Chapter I, Section 9.3], $i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$ is a ${\rm GL}_{n}(E)$ stable $\Lambda$-lattice in $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$–which is also stable under the action of $\Gamma$ on $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$, defined by
@@
Theorem 6.3.1.
Let $\pi$ be a smooth, depth-zero, integral, $\ell$-adic generic representation of ${\rm GL}_{n}(F)$, with $J_{\ell}(\pi)$ the unique generic sub-quotient of the $-\ell$ reduction of $\pi$. Let $\Pi$ be an absolutely irreducible, depth-zero, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. Then the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is the unique generic sub-quotient of the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L}^{\prime})$.
@@
and each $\xi_{i}$ is an irreducible, integral, $\ell$-adic cuspidal representation of ${\rm GL}_{n_{i}}(F)$. Now, $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$, being the base change lift of $\pi$, is a subrepresentation of $i_{P(E)}^{G_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where each $\eta_{i}$ is an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ such that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$. Then we have the following commutative diagram
@@
where $\varphi$ is a ${\rm GL}_{n}(F)$ equivariant map induced from the injection $\mathcal{L^{\prime}}\hookrightarrow i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$. Note that the twisted Jacquet modules $\mathcal{L}^{\prime}_{N_{n}(E),\Theta_{E}}$ and $(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})_{N_{n}(E),\Theta_{E}}$ are free $\Lambda$-modules of rank 111. Also note that the vertical maps $f_{1}$ and $f_{2}$ are non-zero and hence isomorphisms. Therefore, the map $\varphi$ takes the generic subquotients of $\widehat{H}^{0}(\mathcal{L}^{\prime})$ to the generic subquotients of $\widehat{H}^{0}(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})$. Recall that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\xi_{i}$, and $\mathcal{L}_{i}$ is a ${\rm GL}_{n_{i}}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $\eta_{i}$, for each $1\leq i\leq r$. Then using Theorem 4.1.1, we get that the Tate cohomology group $\widehat{H}^{0}(\mathcal{L}_{i})$ is isomorphic to $r_{\ell}(\xi_{i})^{(\ell)}$, and the equation (2.1) gives the isomorphism
@@
Thus, the supercuspidal support of the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is same as the supercuspidal support of the generic subquotients of the Tate cohomology $\widehat{H}^{0}(\mathcal{L}^{\prime})$. Hence, $J_{\ell}(\pi)^{(\ell)}$ is a subquotient of $\widehat{H}^{0}(\mathcal{L}^{\prime})$. This completes the proof. ∎